OpenStudy (anonymous):

what is the common difference of a 43-term arithmetic sequence where the first term is -13 and the sum is 9374

OpenStudy (sidsiddhartha):

first find the last term using \[n(first term+last term)/2=summation\]

OpenStudy (anonymous):

There is no given last term, How would I find it?

OpenStudy (sidsiddhartha):

that is what u have to find n=43 first term=-13 summation=9374 now put them in the formula and find the last term

OpenStudy (anonymous):

43(-13+lastterm)/2=9374

OpenStudy (sidsiddhartha):

yeah :)

OpenStudy (anonymous):

So then would I multiply -13 and (x) by 43 to give me (-559+ 43(x) )/2= 9347

OpenStudy (sidsiddhartha):

\[\frac{ 43(-13+last term) }{ 2 }=9374\] \[-13+lasterm=\frac{ 9374*2 }{43 }\] \[last term=436+13=449\] ok now use this formula \[last term=first term+(n-1)*(common ratio)\] your last term=449 first term=-13 n=43 so \[449=-13+(43-1)*common ratio=-13+42*(common ratio)\] common ratio=11 do u get it??

OpenStudy (anonymous):

Yes, Thankyou.

OpenStudy (sidsiddhartha):

np:)